If the supply equation is given as Q=-200+10P, the price elasticity of supply between R40 and R50, calculated using the arc elasticity of demand, is?
1
Expert's answer
2018-01-28T09:16:08-0500
Arc elasticity of demand is ((Q2 – Q1)/midpoint Q)/((P2 – P1)/midpoint P)
Q1=-200+10P = 40, P=24
Q2=-200+10P = 50, P=25
Arc elasticity of demand = ((50-40)/45)/((25-24)/24,5)=5.44